Which of the following comes closest to the internal rate of return (IRR) of a project that requires an initial investment of $100 and produces a single cash flow of $160 at the end of year 11? The required rate of return for the project is 13%. a. 5.36% b. 4.81 % C. 3.99% d. 4.37% e. 6.05% Oc.
Answer #1

Answer is d. 4.37%

The above can be arrived as

IRR = (Expected Cash flow/ Initial Outlay)^(1/Number of Periods)-1

=(160/100)^(1/11)-1

=(1.6^.0909)-1

=1.0436-1= 0.0436 or 4.36%

Thus, Internal Rate of Return (IRR) is 4.37%